# Non-local Hamilton–Jacobi-Bellman equations for the stochastic optimal control of path-dependent piecewise deterministic processes

### v1

@@ 5.1. The operator $G$
(i) Existence: By Lemma 5.7, $\bar{\psi}=\bar{\psi}(s,x)$ is Lipschitz continuous in $x$. By Theorem 5.4, $\bar{\psi}\in\mathrm{BLSA}([0,T]\times\Omega)$ and
$\bar{\psi}=G\bar{\psi}=G^{2}\bar{\psi}.$
Thus, by Theorem 5.9 (a), $\bar{\psi}$ is a minimax solution of ($\text{{TVP}}_{\psi}$) with $\psi=\bar{\psi}$.
@@
(ii) Uniqueness: Let $u$ be a minimax solution of (TVP). Then $u$ is also a minimax solution of ($\text{{TVP}}_{\psi}$) with $\psi=u$. By Theorem 5.9 (b), $Gu$ is the unique minimax solution of ($\text{{TVP}}_{\psi}$) with $\psi=u$. Moreover, by Theorem 7.7 in [BK1],
$u=Gu.$
This yields $u=\bar{\psi}$ according to Theorem 5.4. ∎
Proof of Theorem 5.2. It remains to note that Theorem 5.2 is an immediate consequence of Theorems 5.4 and 5.11. ∎
@@ 5.3. The comparison principle
Let $u_{0}$ be a minimax sub- and $v_{0}$ be a minimax supersolution of (TVP). Then $u_{0}\leq v_{0}$.
@@
for each $n\in{\mathbb{N}}$. As $u_{0}$ and $v_{0}$ are semicontinuous, we can with the same argument as in Theorem 5.9 (b) deduce that $u_{N,1}$ (resp. $v_{N,1}$) is the unique minimax solution of ($\text{{TVP}}_{\psi}$) on $[r_{N-1},r_{N}]\times\Omega$ with $\psi=u_{0}$ (resp., $\psi=v_{0}$).
Note that $u_{0}$ (resp., $v_{0}$) is also a minimax subsolution (resp., supersolution) to ($\text{{TVP}}_{\psi}$) with $\psi=u_{0}$. Thus, by the comparison principle for ($\text{{TVP}}_{\psi}$) (Corollary  5.4 in [BK1]), $u_{0}\leq u_{N,1}$ and $v_{N,1}\leq v_{0}$ on $[r_{N-1},r_{N}]\times\Omega$.
By induction, one can similarly show that $(u_{N,n})$ is non-decreasing and that $(v_{N,n})$ is non-increasing. Moreover, by Theorem 5.4 (ii), $u_{N,n}\to\bar{\psi}$ and $v_{N,n}\to\bar{\psi}$ uniformly on $[r_{N-1},r_{N}]\times\Omega$ as $n\to\infty$. Thus $u_{0}\leq\bar{\psi}\leq v_{0}$ on $[r_{N-1},r_{N}]\times\Omega$.
Step 2. Let us next consider the interval $[r_{N-2},r_{N}]$. We can proceed as in Step 1 and arrive at the corresponding conclusion but need to replace $h$ by $\bar{\psi}$.
Step 3. Iterating the steps above yields $u_{0}\leq\bar{\psi}\leq v_{0}$ on $[0,T]\times\Omega$. ∎
@@ Appendix A A general path-dependent discrete-time infinite-horizon decision model
We set up a discrete-time infinite-horizon decision model with data that depend on past states and past controls. Then we formulate a corresponding discrete-time infinite-horizon stationary decision model (see also chapter 10 of [BS] regarding a related procedure concerning non-stationary models, but note that our situation is slightly more involved). Finally, we establish part of the Bellman equation for the former model by using the latter model.
@@ A.1. A model with data that depend on past states and past controls
Our goal is to minimize a certain expected cost over the space $\Pi^{\prime}$ of all policies $\pi=(\mu_{0},\mu_{1},\ldots)$, where each
@@
We use randomized policies for two reasons. First, by doing so, we operate immediately in the setting of [BS], whose results we need to use. Second, the notation for non-randomized policies would become very cumbersome in our model. E.g., (realized) policies would then be of the form $\mu_{0}(x_{0})$, $\mu_{1}(x_{0},x_{1};\mu_{0}(x_{0}))$, $\mu_{2}(x_{0},x_{1},x_{2};\mu_{0}(x_{0}),\mu_{1}(x_{0},x_{1};\mu_{0}(x_{0})))$, and so on. The randomized counterparts are of the form $\mu_{0}({\mathrm{d}}u_{0}|x_{0})$, $\mu_{1}({\mathrm{d}}u_{1}|x_{0},x_{1};u_{0})\,\mu_{0}({\mathrm{d}}u_{0}|x_{0})$, $\ldots$,
@@
(A.4) $\begin{split}\int h_{k}\,{\mathrm{d}}r^{k}_{k+1}(\pi^{k},p^{k})&:=\int_{S^{k+1}\times U^{k}}\int_{U}h_{k}(x_{0},\ldots,x_{k};u_{0},\ldots,u_{k})\,\\ &\qquad\qquad\mu_{k}({\mathrm{d}}u_{k}|x_{0},\ldots,x_{k};u_{0},\ldots,u_{k-1})\\ &\qquad\qquad p^{k}({\mathrm{d}}x_{0}\,\ldots\,{\mathrm{d}}x_{k}\,{\mathrm{d}}u_{0},\ldots\,{\mathrm{d}}u_{k-1}),\\ \int h_{N}\,{\mathrm{d}}r^{k}_{N+1}(\pi^{k},p^{k})&:=\int\Big{[}\int_{S}\int_{U}h_{N}(x_{0},\ldots,x_{N};u_{0},\ldots,u_{N})\\ &\qquad\qquad\mu_{N}({\mathrm{d}}u_{N}|x_{0},\ldots,x_{N};u_{0},\ldots,u_{N-1})\\ &\qquad\qquad p_{N-1}({\mathrm{d}}x_{N}|x_{0},\ldots,x_{N-1};u_{0},\ldots,u_{N-1})\Big{]}\\ &\,r^{k}_{N}(\pi^{k},p^{k})({\mathrm{d}}x_{0}\,{\mathrm{d}}u_{0}\,\ldots\,{\mathrm{d}}x_{N-1}\,{\mathrm{d}}u_{N-1})\end{split}$
@@
Moreover, we show that minimization over non-randomized policies is equivalent to minimization over randomized policies. We want to emphasize that we still work with the data introduced at the beginning of this section (section A.1). In particular, non-negativity of the cost functions is important. This will enable us to apply corresponding results in [BS].
@@
Let $x\in S$. Let $J^{\ast}$ and $J_{\pi}$ be the functions defined respectively by (A.3) and (A.2) Then
(A.8) $J^{\ast}(x)=\inf_{\pi\in\Pi}J_{\pi}(x),$
where $\Pi$ is the set of all non-randomized policies $\pi=(\mu_{0},\mu_{1},\ldots)\in\Pi^{\prime}$, i.e., for every $k\in{\mathbb{N}}_{0}$, every $x_{0}$, $\ldots$, $x_{k}\in S$, and every $u_{0}$, $\ldots$, $u_{k-1}\in U$, the measure $\mu_{k}(du_{k}|x_{0},\ldots,x_{k};u_{0},\ldots u_{k-1})$ is a Dirac measure.
@@ A.2. A corresponding stationary model
  • State space $\mathbf{S}:={\mathbb{N}}_{0}\times S^{{\mathbb{N}}_{0}}\times U^{{\mathbb{N}}_{0}}$. A typical state is denoted by
@@
  • Control space $U$.
  • Disturbance space $S$.
@@
    $\mathbf{p}({\mathrm{d}}w|\mathbf{x},u):=p_{k}({\mathrm{d}}w|x^{0},\ldots,x^{k};u^{0},\ldots,u^{k-1},u)$
    on $\mathbf{S}$ given $\mathbf{S}\times U$.
@@
    $\mathbf{f}(\mathbf{x},u,w):=(k+1;x^{0},\ldots,x^{k},w,x^{k+2},x^{k+3},\ldots;u^{0},\ldots,u^{k-1},u,u^{k+1},u^{k+2},\ldots).$
@@
    $\mathbf{g}(\mathbf{x},u):=g_{k}(x^{0},\ldots,x^{k};u^{0},\ldots,u^{k-1},u).$
@@
The set of all Markov policies, denoted by $\vec{\Pi}$, consists of all $\vec{\pi}=(\vec{\mu}_{0},\vec{\mu}_{1},\ldots)\in\vec{\Pi}^{\prime}$, where each $\vec{\mu}_{k}$, $k\in{\mathbb{N}}_{0}$, only depends on $\mathbf{x}_{k}$ instead of $\mathbf{x}_{0}$, $\ldots$, $\mathbf{x}_{k}$, $u_{0}$, $\ldots$, $u_{k-1}$, and in this case we write
$\vec{\mu}_{k}({\mathrm{d}}u_{k}|\mathbf{x}_{k})$
instead of $\vec{\mu}_{k}({\mathrm{d}}u_{k}|\mathbf{x}_{0},\,\ldots,\,\mathbf{x}_{k};\,u_{0},\,\ldots,u_{k-1})$.
@@
which we will do by proving the following four claims. Note that we will use the following notational conventions for the components of elements $(\mathbf{x}_{0},\mathbf{x}_{1},\ldots)$ of $\mathbf{S}^{{\mathbb{N}}_{0}}$:
(A.14) $\mathbf{x}_{k}=(\iota_{k};x^{0}_{k},x^{1}_{k},\ldots;u^{0}_{k},u^{1}_{k},\ldots),\quad k\in{\mathbb{N}}_{0}.$
@@
$\qquad\cdot p_{1}({\mathrm{d}}x^{2}_{1}|x,w;u,u_{0})\,\prod_{j=3}^{\infty}\delta_{x}({\mathrm{d}}x^{j}_{1})\,\delta_{u}({\mathrm{d}}u^{0}_{1})\,\delta_{u_{0}}({\mathrm{d}}u^{1}_{1})\,\prod_{j=2}^{\infty}\delta_{u_{\Delta}}({\mathrm{d}}u^{j}_{1}),\text{ etc.}$.
@@
Proof of Claim 3. Fix a policy $\pi=(\mu_{0},\mu_{1},\ldots)\in\Pi^{\prime}$. Define a policy $\vec{\pi}=(\vec{\pi}_{0},\vec{\pi}_{1},\ldots)\in\vec{\Pi}^{\prime}$ by
@@
$\vec{\pi}_{k}({\mathrm{d}}u_{k}|\mathbf{x}_{0},\ldots,\mathbf{x}_{k};u_{0},\ldots,u_{k-1}):=\mu_{k+1}({\mathrm{d}}u_{k}|x,w,x^{2}_{1},\ldots,x^{k+1}_{k};u,u_{0},\ldots,u_{k-1})$
@@
Next, following the proof of Claim 1 in section A.3, we can deduce via defining a corresponding non-randomized policy in $\Pi$ (see Theorem A.4 for its definition) for any non-randomized policy in $\vec{\Pi}^{\prime}$ that
@@ Appendix B Some technical proofs
Consider the maps $\iota_{i}(s,x)\mapsto\iota_{i}(s,x)$, +×${\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\to\Omega$, $i\in\{1,2\}$, and $\iota_{3}:(s,x,\alpha)\mapsto\iota_{3}(s,x,\alpha)$, +×${\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\to{\mathbb{R}}_{+}\times\Omega$, respectively defined by
@@
where $\iota_{2}^{t}$ is defined by (B.1) below. We have that
@@
Proof of Lemma B.1. Item (i) directly follows from (ii)-(iii), noticing that $\iota_{1}=\iota_{2}\circ\iota_{3}$. Item (iii) is an adaption of Theorem 96 (b), 146-IV, in [DMprobPot]. In order to prove item (ii), we consider the map
The proof follows the line of the reasoning on pp. 84 and 135 in [Billingsley] regarding the measurability of path-space valued random elements, using that
@@
which we show next. First note that the maps $s\mapsto\int_{0}^{t}f(s+r,x,\alpha(s+r))\,{\mathrm{d}}r=\int_{s}^{s+t}f(r,x,\alpha(r))\,{\mathrm{d}}r$, ${\mathbb{R}}_{+}\to{\mathbb{R}}^{d}$, $(x,\alpha)\in\Omega\times\mathcal{A}$, are continuous. Moreover, for every $s\in{\mathbb{R}}_{+}$, the map $(x,\alpha)\mapsto\int_{0}^{t}f(s+r,x,\alpha(s+r))\,{\mathrm{d}}r$, $\Omega\times\mathcal{A}\to{\mathbb{R}}^{d}$, is ${\mathcal{F}}^{0}_{t+s}\otimes\mathcal{B}(\mathcal{A})$- and hence ${\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$-measurable.171717 To see this, consider the case that each component $f^{j}$ of $f=(f^{1},\ldots,f^{d})$ is of the form $f^{j}(r,x,a)=f^{j}_{1}(x(\cdot\wedge(t+s)))\,f^{j}_{2}(r,a)$, where $f^{j}_{1}$ is ${\mathcal{F}}^{0}$- and $f^{j}_{2}$ is (+)×$\mathcal{B}({\mathbb{R}}_{+})\times\mathcal{B}(A)$-measurable. Taking Remark 3.1 into account, we can deduce the ${\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$-measurability of $(x,\alpha)\mapsto\int_{0}^{t}f(s+r,x,\alpha(s+r))\,{\mathrm{d}}r$. It remains to apply a monotone-class argument. The reasoning here is very similar to the proof of Lemma 2 in [Yushkevich80]. These two facts establish (ii’) (see  Proposition 1.13, p. 5, in [KaratzasShreve]). ∎
@@
The map $(s,x,\alpha)\mapsto\phi^{s,x,\alpha}_{(0)}$, +×${\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\to\Omega$, defined by
@@
Proof of Lemma B.2. By Theorem 96 (d), 146-IV, in [DMprobPot], the map $\iota_{4}:(s,x,\tilde{x})\mapsto\iota_{4}(s,x)$, +×${\mathbb{R}}_{+}\times\Omega\times\Omega\to\Omega$, defined by
@@
is measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes{\mathcal{F}}^{0}$ to ${\mathcal{F}}^{0}$. This allows to conclude the proof, being $\phi^{s,x,\alpha}_{(0)}=\iota_{4}(s,x,\iota_{1}(s,x,\alpha))$, and recalling by Lemma B.1-(i) the measurability of $(s,x,\alpha)\mapsto(s,x,\iota_{1}(s,x,\alpha))$, +×${\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\to{\mathbb{R}}_{+}\times\Omega$, from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes{\mathcal{F}}^{0}$. ∎
@@
Consider the mappings $(s,x,\alpha)\mapsto\phi^{s,x,\alpha}_{(n)}$, +×${\mathbb{R}}_{+}\times\Omega\to\Omega$, $n\in{\mathbb{N}}$, recursively defined by

### v2

@@ 5.4. Existence and uniqueness
(i) Existence: By Theorem 5.12, $\bar{\psi}=\bar{\psi}(s,x)$ is Lipschitz continuous in $x$. By Theorem 5.10, $\bar{\psi}\in\mathrm{BLSA}([0,T]\times\Omega)$ and $\bar{\psi}=G\bar{\psi}=G^{2}\bar{\psi}$. Thus, by Theorem 5.8 (a), $\bar{\psi}$ is a minimax solution of ($\text{{TVP}}_{\psi}$) with $\psi=\bar{\psi}$. We can conclude that $\bar{\psi}$ is a minimax solution of (5.1).
@@
(ii) Uniqueness: Let $u$ be a minimax solution of (5.1). Then $u$ is also a minimax solution of ($\text{{TVP}}_{\psi}$) with $\psi=u$. By Theorem 5.8 (b), GuGu is the unique minimax solution of ($\text{{TVP}}_{\psi}$) with $\psi=u$. Moreover, by Theorem 7.7 in [BK1], $u=Gu$, which yields $u=\bar{\psi}$ according to Theorem 5.10 (i). This concludes the proof. ∎
The next result is an immediate consequence of Theorems  5.13 and  5.10.
Corollary 5.14.
The value function $V$ from (4.6) is the unique minimax solution of (5.1) on $(0,T]\times\Omega$.
@@ 5.5. The comparison principle
Let $u_{0}$ be a minimax sub- and $v_{0}$ be a minimax supersolution of (5.1). Then $u_{0}\leq v_{0}$.
@@
for each $n\in{\mathbb{N}}$. As $u_{0}$ and $v_{0}$ are semicontinuous, we can with the same argument as in Theorem 5.8 (b) deduce that $u_{N,1}$ (resp. $v_{N,1}$) is the unique minimax solution of ($\text{{TVP}}_{\psi}$) on $r_{N-1},r_{N}]\times\Omega$ with $\psi=u_{0}$ (resp., $\psi=v_{0}$).
Note that $u_{0}$ (resp., $v_{0}$) is also a minimax subsolution (resp., supersolution) to ($\text{{TVP}}_{\psi}$) with $\psi=u_{0}$. Thus, by the comparison principle for ($\text{{TVP}}_{\psi}$) (Corollary  5.4 in [BK1]), $u_{0}\leq u_{N,1}$ and $v_{N,1}\leq v_{0}$ on $r_{N-1},r_{N}]\times\Omega$.
By induction, one can similarly show that $(u_{N,n})$ is non-decreasing and that $(v_{N,n})$ is non-increasing. Moreover, by Theorem 5.10 (ii), $u_{N,n}\to\bar{\psi}$ and $v_{N,n}\to\bar{\psi}$ uniformly on $r_{N-1},r_{N}]\times\Omega$ as $n\to\infty$. Thus $u_{0}\leq\bar{\psi}\leq v_{0}$ on $r_{N-1},r_{N}]\times\Omega$.
Step 2. Let us next consider the interval $r_{N-2},r_{N}]$. We can proceed as in Step 1 and arrive at the corresponding conclusion but need to replace $h$ by $\bar{\psi}$.
Step 3. Iterating the steps above yields $u_{0}\leq\bar{\psi}\leq v_{0}$ on $0,T]\times\Omega$. ∎
@@ Appendix A A general path-dependent discrete-time infinite-horizon decision model
We set up a discrete-time infinite-horizon decision model with data that depend on past states and past controls. Then we formulate a corresponding discrete-time infinite-horizon stationary decision model (see also chapter 10 of [BS] regarding a related procedure concerning non-stationary models). Finally, we establish part of the Bellman equation for the former model by using the latter model. Note that here the terms stationary and non-stationary have the same meaning as in [BS].
@@ A.1. A model with data that depend on past states and past controls
Our goal is to minimize a certain expected cost over the space $\Pi^{\prime}$ of all randomized policies $\pi=(\mu_{0},\mu_{1},\ldots)$, where each
@@
Later we will also use the space $\Pi$ of all non-randomized policies $\pi=(\mu_{0},\mu_{1},\ldots)\in\Pi^{\prime}$, i.e., for every $k\in{\mathbb{N}}_{0}$, every $x_{0}$, $\ldots$, $x_{k}\in S$, and every $u_{0}$, $\ldots$, $u_{k-1}\in U$, the measure $\mu_{k}(du_{k}|x_{0},\ldots,x_{k};u_{0},\ldots u_{k-1})$ is a Dirac measure.
@@
We use randomized policies for two reasons. First, by doing so, we operate immediately in the setting of [BS], whose results we need to use. Second, the natural notation for non-randomized policies would become very cumbersome in our model. E.g., (realized) policies would then be of the form $\mu_{0}(x_{0})$, $\mu_{1}(x_{0},x_{1};\mu_{0}(x_{0}))$, $\mu_{2}(x_{0},x_{1},x_{2};\mu_{0}(x_{0}),\mu_{1}(x_{0},x_{1};\mu_{0}(x_{0})))$, and so on. Using the notation for randomized policies, we have instead $\mu_{0}({\mathrm{d}}u_{0}|x_{0})$, $\mu_{1}({\mathrm{d}}u_{1}|x_{0},x_{1};u_{0})\,\mu_{0}({\mathrm{d}}u_{0}|x_{0})$, $\ldots$,
@@
(A.4) $\begin{split}\int h_{k}\,{\mathrm{d}}r^{k}_{k+1}(\pi^{k},p^{k})&:=\int_{S^{k+1}\times U^{k}}\int_{U}h_{k}(x_{0},\ldots,x_{k};u_{0},\ldots,u_{k})\,\\ &\qquad\qquad\mu_{k}({\mathrm{d}}u_{k}|x_{0},\ldots,x_{k};u_{0},\ldots,u_{k-1})\\ &\qquad\qquad p^{k}({\mathrm{d}}x_{0}\,\ldots\,{\mathrm{d}}x_{k}\,{\color[rgb]{0,0,0}{\mathrm{d}}u_{0}\ldots\,{\mathrm{d}}u_{k-1}}),\\ \int h_{N}\,{\mathrm{d}}r^{k}_{N+1}(\pi^{k},p^{k})&:=\int\Big[\int_{S}\int_{U}h_{N}(x_{0},\ldots,x_{N};u_{0},\ldots,u_{N})\\ &\qquad\qquad\mu_{N}({\mathrm{d}}u_{N}|x_{0},\ldots,x_{N};u_{0},\ldots,u_{N-1})\\ &\qquad\qquad p_{N-1}({\mathrm{d}}x_{N}|x_{0},\ldots,x_{N-1};u_{0},\ldots,u_{N-1})\Big]\\ &\,r^{k}_{N}(\pi^{k},p^{k})({\mathrm{d}}x_{0}\,{\mathrm{d}}u_{0}\,\ldots\,{\mathrm{d}}x_{N-1}\,{\mathrm{d}}u_{N-1})\end{split}$
@@
Moreover, we show that minimization over non-randomized policies, i.e., over $\Pi$, is equivalent to minimization over randomized policies, i.e., over $\Pi^{\prime}$. We want to emphasize that we still work with the data introduced at the beginning of this section (section A.1). In particular, non-negativity of the cost functions is important. This will enable us to apply corresponding results in [BS].
@@
Let $x\in S$. Let $J^{\ast}$ and $J_{\pi}$ be the functions defined respectively by (A.3) and (A.2). Then
(A.8) $J^{\ast}(x)=\inf_{\pi\in\Pi}J_{\pi}(x).$
@@ A.2. A corresponding stationary model
  • State space $\mathbf{S}:={\mathbb{N}}_{0}\times S^{{\mathbb{N}}_{0}}\times U^{{\mathbb{N}}_{0}}$, which is a non-empty Borel space (see, e.g., Proposition 7.13 in [BS]). A typical state is denoted by
@@
  • Control space $U$, which is a non-empty Borel space.
  • Disturbance space $S$, which is a non-empty Borel space.
@@
    $\mathbf{p}({\mathrm{d}}w|\mathbf{x},u):=p_{k}({\mathrm{d}}w|x^{0},\ldots,x^{k};u^{0},\ldots,u^{k-1},u),$
    which is a Borel-measurable stochastic kernel on $S$ given $\mathbf{S}\times U$.
@@
    $\mathbf{f}(\mathbf{x},u,w):=(k+1;x^{0},\ldots,x^{k},w,x^{k+2},x^{k+3},\ldots;u^{0},\ldots,u^{k-1},u,u^{k+1},u^{k+2},\ldots),$
    which is a Borel measurable function.
@@
    $\mathbf{g}(\mathbf{x},u):=g_{k}(x^{0},\ldots,x^{k};u^{0},\ldots,u^{k-1},u),$
    which is a lower semi-analytic function.
Remark A.5.
Note that our data satisfy all assumptions of the infinite-horizon model in Chapter 9 of [BS]. This allows us to to apply the results of [BS] later in sections A.3 and A.4.
@@
Remark A.6.
As an alternative to the stationary model introduced above, one could consider instead of our state space $\mathbf{S}={\mathbb{N}}_{0}\times S^{{\mathbb{N}}_{0}}\times U^{{\mathbb{N}}_{0}}$ the set
$\bigcup_{k\in{\mathbb{N}}_{0}}\{(k;x^{0},\ldots,x^{k};u^{0},\ldots,u^{k-1}):\,(x^{0},\ldots,x^{k})\in S^{k+1},\,(u^{0},\ldots,u^{k-1})\in U^{k}\}$
as state space and then appropriately adjust the remaining parts of the model such as the disturbance kernel, the system function, and so on. Doing so would be more in line with section 10.1 of [BS], where one can find a similar procedure of transforming a non-stationary to a stationary model.
@@
which we will do by proving the following four claims.
@@
Note that in the proofs of these claims, we will use the following notational conventions for the components of elements $(\mathbf{x}_{0},\mathbf{x}_{1},\ldots)$ of $\mathbf{S}^{{\mathbb{N}}_{0}}$:
(A.14) $\mathbf{x}_{k}=(\iota_{k};x^{0}_{k},x^{1}_{k},\ldots;u^{0}_{k},u^{1}_{k},\ldots),\quad k\in{\mathbb{N}}_{0},$
where $\iota_{k}\in{\mathbb{N}}_{0}$ denotes the first component of $\mathbf{x}_{k}$, $x^{0}_{k}\in S$ the second one, and so on.
@@
$\qquad\cdot p_{1}({\mathrm{d}}x^{2}_{1}|x,w;u,u_{0})\,\prod_{j=3}^{\infty}\delta_{x}({\mathrm{d}}x^{j}_{1})\,\delta_{u}({\mathrm{d}}u^{0}_{1})\,\delta_{u_{0}}({\mathrm{d}}u^{1}_{1})\,\prod_{j=2}^{\infty}\delta_{u_{\Delta}}({\mathrm{d}}u^{j}_{1}),\text{ etc.}$
@@
Proof of Claim 3. Fix a policy $\pi=(\mu_{0},\mu_{1},\ldots)\in\Pi^{\prime}$. Define a policy $\vec{\pi}={\color[rgb]{0,0,0}(\vec{\mu}_{0},\vec{\mu}_{1},\ldots)}\in\vec{\Pi}^{\prime}$ by
@@
${\color[rgb]{0,0,0}\vec{\nu}_{k}}({\mathrm{d}}u_{k}|\mathbf{x}_{0},\ldots,\mathbf{x}_{k};u_{0},\ldots,u_{k-1}):=\mu_{k+1}({\mathrm{d}}u_{k}|x,w,x^{2}_{1},\ldots,x^{k+1}_{k};u,u_{0},\ldots,u_{k-1})$
@@
Next, following the proof of Claim 1 in section A.3, we can deduce via defining a corresponding non-randomized policy in $\Pi$ for any non-randomized policy in $\vec{\Pi}^{\prime}$ that
@@ Appendix B Some technical proofs
Consider the maps $(s,x,{\color[rgb]{0,0,0}\alpha})\mapsto\iota_{i}(s,x,{\color[rgb]{0,0,0}\alpha})$, ${\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\to\Omega$, $i\in\{1,2\}$, and $\iota_{3}:(s,x,\alpha)\mapsto\iota_{3}(s,x,\alpha)$, ${\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\to{\mathbb{R}}_{+}\times\Omega{\color[rgb]{0,0,0}\times\mathcal{A}}$, respectively defined by
@@
where $\iota_{2}^{t}$ is given by
We have that
@@
Proof of Lemma B.1. Item (i) directly follows from (ii)-(iii), noticing that $\iota_{1}=\iota_{2}\circ\iota_{3}$. Item (iii) follows from the measurability of $(s,x,\alpha)\mapsto x(\cdot\wedge s)$, ${\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\to\Omega$, from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to ${\mathcal{F}}^{0}$ due to Theorem 96 (b), 146-IV, in [DMprobPot] together with Proposition 7.14 in [BS]. In order to prove item (ii), we consider the map in (B.1). The proof follows the line of the reasoning on p. 135 in [Billingsley] regarding the measurability of path-space valued random elements, using that
@@
which we show next.
First note that the maps
$s\mapsto\int_{0}^{t}f(s+r,x,\alpha(s+r))\,{\mathrm{d}}r=\int_{s}^{s+t}f(r,x,\alpha(r))\,{\mathrm{d}}r,\,{\mathbb{R}}_{+}\to{\mathbb{R}}^{d},\,(x,\alpha)\in\Omega\times\mathcal{A},$
are continuous.
Moreover, for every $s\in{\mathbb{R}}_{+}$, the map
$(x,\alpha)\mapsto\int_{0}^{t}f(s+r,x,\alpha(s+r))\,{\mathrm{d}}r,\,\Omega\times\mathcal{A}\to{\mathbb{R}}^{d},$
is ${\mathcal{F}}^{0}_{t+s}\otimes\mathcal{B}(\mathcal{A})$- and hence ${\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$-measurable. To see this, consider the case that each component $f^{j}$ of $f=(f^{1},\ldots,f^{d})$ is of the form
$f^{j}(r,x,a)=f^{j}_{1}(x(\cdot\wedge(t+s)))\,f^{j}_{2}(r,a),$
where $f^{j}_{1}$ is ${\mathcal{F}}^{0}$- and $f^{j}_{2}$ is $\mathcal{B}({\mathbb{R}}_{+})\times\mathcal{B}(A)$-measurable. Taking Remark 3.1 into account, we can deduce the ${\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$-measurability of
$(x,\alpha)\mapsto\int_{0}^{t}f(s+r,x,\alpha(s+r))\,{\mathrm{d}}r.$
It remains to apply a monotone-class argument. The reasoning here is very similar to the proof of Lemma 2 in [Yushkevich80].
The above two facts establish (ii’) (see  Proposition 1.13, p. 5, in [KaratzasShreve]). ∎
@@
The map $(s,x,\alpha)\mapsto\phi^{s,x,\alpha}_{(0)}$, ${\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\to\Omega$, defined by
@@
Proof of Lemma B.2. By Theorem 96 (d), 146-IV, in [DMprobPot], the map $\iota_{4}:(s,x,\tilde{x})\mapsto\iota_{4}(s,x)$, ${\mathbb{R}}_{+}\times\Omega\times\Omega\to\Omega$, defined by
@@
is measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes{\mathcal{F}}^{0}$ to ${\mathcal{F}}^{0}$. This allows to conclude the proof, being $\phi^{s,x,\alpha}_{(0)}=\iota_{4}(s,x,\iota_{1}(s,x,\alpha))$, and recalling by Lemma B.1-(i) the measurability of $(s,x,\alpha)\mapsto(s,x,\iota_{1}(s,x,\alpha))$, ${\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\to{\mathbb{R}}_{+}\times\Omega$, from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes{\mathcal{F}}^{0}$. ∎
@@
Consider the mappings $(s,x,\alpha)\mapsto\phi^{s,x,\alpha}_{(n)}$, ${\mathbb{R}}_{+}\times\Omega\to\Omega$, $n\in{\mathbb{N}}$, recursively defined by